Solve the system using substitution.
3x + 4y − z = 14
x − 5y + 2z = 26
5x + y − 2z = 40

The solution to the system of three linear equations is the ordered triple____.

Answer by Alan3354(69443)

Solve the system using substitution.
3x + 4y - z = 14 times 2 --> 6x + 8y - 2z = 28
x - 5y + 2z = 26
5x + y - 2z = 40
-------
6x + 8y - 2z = 28 --> 2z = 6x + 8y - 28
x - 5y + 2z = 26
5x + y - 2z = 40
==========
Sub for 2z in Eqns 2 & 3
x - 5y + 2z = 26
x - 5y + 6x + 8y - 28 = 26
7x + 3y = 54 --- Eqn A
---
5x + y - 2z = 40
5x + y - (6x + 8y - 28) = 40
-x - 7y = 12 --- Eqn B
-7x - 49y = 84 --- Eqn B times 7
7x - y = 54 --- Eqn A
-------------------------------- Add
-50y = 138
y = -69/25
============
Sub for y in Eqn A or B, find x
Sub for x & y in any equation, find z
========================
Don't solve for x, y or z and deal with fractions.
------
check the work carefully.
It's easy to make a mistake, and it's late and I'm tired.